on three consecutive passws, a football teams gains 6 yards, loses 26 yards, and gains 23 yards. what number represents the total net yardage?
Answer by stanbon(75887)

on three consecutive passes, a football teams gains 6 yards, loses 26 yards, and gains 23 yards. what number represents the total net yardage?
-----
6 - 26 + 23 = 29-26 = 3 yds gained
=======================
Cheers,
Stan H.